OpenStudy (anonymous):

he amount of a radioactive material changes with time. The table below shows the amount of radioactive material f(t) left after time t: t(hours) 0 1 2 f(t) 180 90 45 Which exponential function best represents the relationship between f(t) and t? f(t) = 0.5(180)t

OpenStudy (anonymous):

@cwrw238 @ganeshie8

OpenStudy (aum):

You have provided only one answer choice and that is not the answer. Try, t = 0, t = 1 and t = 2 in each choice and see if they output 180, 90 and 45 respectively.

OpenStudy (anonymous):

(t) = 180(0.5)t

OpenStudy (aum):

f(t) = 180(0.5)^t put t = 0 What do you get for f(t) ?

OpenStudy (anonymous):

0.5

OpenStudy (anonymous):

0.5*180=90

OpenStudy (anonymous):

0.5*90=45

OpenStudy (aum):

When you put t = 0 in 180(0.5)^t you get 180 * (0.5)^0 = 180 * 1 = 180 t = 1, f(t) = 180 * 0.5 = 90 t = 2, f(t) = 180 * .5 * .5 = 45 So it satisfies all three f(t) values in the table.

OpenStudy (anonymous):

so i was right